# Constrained 6-DoF Grasp Generation on Complex Shapes for Improved Dual-Arm Manipulation

### v1

@@ IV CGDF: Constrained Grasp Diffusion Fields
Given a object point cloud $P\in\mathbb{R}^{3}$ along with a target region (i.e. constraint) $P_{t}\subseteq P$, our goal is to generate $M$ parallel-jaw grasp poses $H_{i}\in SE(3),i\in[0,M)$ on $P$, such that $H_{i}$ is located on $P_{t}$. Here $M$ is a number that can be user-defined. Additionally, we make no assumptions regarding the granularity of the target regions, allowing $P_{t}$ to encompass anything from a small, localized region to the entire point cloud. For an n-arm setting with large objects, $n$ target regions can be selected $P_{T}=\{P_{t}^{1},P_{t}^{2}\dots P_{t}^{n}\}$. Consequently, our method must learn a robust shape representation to generalize to arbitrary shapes and complex geometries, ensuring its applicability across various constraint regions. Further, the model should generate diverse grasps with full coverage over the target region. Diffusion models have shown excellent performance in generating diverse outputs while learning the data distribution effectively [28]. Thus we choose a diffusion-based architecture as our backbone.
@@ IV-A SE(3) Diffusion Model
$\hat{H}=\text{Expmap}\left[\text{Logmap}(H)+\epsilon_{k}\right],\epsilon_{k}\sim\mathcal{N}(0,\sigma_{k}^{2}I)$ (3)
Loss Function: The diffusion model can now be trained to predict the noise given a perturbed sample by minimizing the L1 loss objective between the sampled noise $\epsilon_{k}$ and the predicted vector as follows:
@@
Inverse Diffusion Step: As shown in Fig 2(c), the inverse diffusion step is formulated as an adapted version of the Euclidean Langevin MCMC  [32] as in  [28]:
$H_{k-1}=\text{Expmap}\left(\frac{\alpha_{k}^{2}}{2}s_{\theta}(H_{k},k,P)+\alpha_{k}\epsilon\right)H_{k}$ (5)
where $\epsilon\in\mathbb{R}^{6},\epsilon_{k}\sim\mathcal{N}(0,I)$ , $\alpha_{k}$ refers to a non-negative step-dependent coefficient.
@@ IV-B Constrained Grasp Diffusion
The above diffusion model trains on an unconstrained grasping objective and thus generates grasps covering the entire object pointcloud $P$. However, given that the model learns rich enough shape representations, we can use the same model without retraining to generate constrained grasps in a sample-efficient manner by using the energy values predicted by $E_{\theta}$. We exploit the fact that grasps that are colliding or bad quality have higher energy values as compared to stable grasps.
@@
Given an object pointcloud $P$ and a target region pointcloud $P_{t}$, we first start with generating a grasp pose $H_{t}$ on $P_{t}$ by treating it as a stand-alone pointcloud. If $H_{t}$ is a stable grasp pose, the final energy $e^{\prime\prime}$ (with $k=0$) of $H_{t}$ using Equation 1, should be low. Now, we calculate the energy $e^{\prime}$ of the $H_{t}$, giving $P$ as the context instead of $P_{t}$. Notice that $e^{\prime}$ is low only if $H_{t}$ is a stable grasp on $P$. Thus, $H_{t}$ is a good constrained grasp if both $e^{\prime}$ and $e^{\prime\prime}$ are low. Formally $H_{t}$ is a stable constrained grasp if:
$e^{\prime}<\delta\And e^{\prime\prime}<\delta$ (6)
where $\delta$ is a user-defined threshold and,
$e^{\prime}=E_{\theta}(H_{t},0,P)$ (7a)
$e^{\prime\prime}=E_{\theta}(H_{t},0,P_{t})$ (7b)
@@
$s_{\theta}(H_{k},k,\{P,P_{t}\})=-\frac{D\text{max}(e^{\prime}_{k},e^{\prime\prime}_{k})}{DH_{k}}$
@@
As shown in Equations 9 and 8, by taking the maximum value of the energies, we essentially guide the grasp from a random pose to a stable configuration near the target region. During the inverse diffusion step, if the grasp $H_{k}$ is near the constrained region but collides with the full object or is unstable, the energy $e^{\prime}_{k}$ is higher, which makes $H_{k}$ move to a more stable pose. Conversely, if the grasp is stable but is far from the constrained region, then $e^{\prime\prime}_{k}$ is higher, and the grasp moves closer to the constrained region. Eventually, the grasp moves to a pose where the energies for both $P$ and $P_{t}$ are low, i.e. the grasp is on the constrained region as well as stable. Since this method is heavily dependent on the energy values, it requires that the EBM $E_{\theta}$ learns a shape representation that is expressive enough to distinguish between arbitrary shapes at a fine-grained level.
@@ IV-C Model Architecture
Our feature representation is inspired by  [27] and  [28], which show that joint learning of surface reconstruction and grasping generates rich grasp features. However, these methods only use a global point cloud embedding as the conditional vector. This works for smaller objects where modeling the intricate geometries are irrelevant for grasp generation but fails to give good results when working with larger objects with complex shapes, as this setting involves modeling local geometries effectively. As shown in Fig 2, we build on  [28] to incorporate local geometric features inspired from  [12]. CGDF first uses a VN-Pointnet  [33] combined with a convolutional plane encoder to generate three canonical feature planes. Given a query point $x_{q}$, the feature encoder $F_{\theta}$ is an MLP that outputs the sdf along with the feature vector
@@
where $\oplus$ means concatenation, and $z_{x_{q}}$ is the feature vector at $x_{q}$ sampled from the feature planes using bilinear interpolation. The grasp pose $H\in SE(3)$ is represented as $P_{H}=H\cdot P_{q}$, where $P_{q}\in\mathbb{R}^{3}$ is a fixed query pointcloud. The pose descriptors of $H$ can then be obtained as $f_{H}^{k}=\bigoplus_{i=0}^{N}[f_{x_{i}}^{k}],x_{i}\in P_{H}$. The decoder is an MLP that takes the grasp pose descriptors and outputs the energy of the grasp, $e_{k}=D_{\theta}(f_{H}^{k})$. The model trains on the main objective as defined in Equation 4, along with an auxiliary loss to predict the SDF as mentioned in [28].
@@ V Experiments
[…]
@@
We evaluate the performance of these models using three metrics: Force Closure, Grasp Success Rate and Target Grasps. Force Closure serves as a stringent criterion, accounting for both stability and collisions between the object and the gripper in the generated dual-arm grasps. On the other hand, the Grasp Success Rate reflects a more practical assessment, simulating scenarios typical of physical interactions within an Isaac Gym environment. Target Grasps measures the percentage of grasps that are generated near the target region.
@@
$GG^{\prime}\succcurlyeq\epsilon I_{6\times 6};\text{ and }\|Gc\|<\delta$ (10)
where $G$ is the grasp matrix, $c$ is the axis of the friction cone, $S$ is the object surface and $x_{i}\in S$. Please refer to  [37] for more details. To get the contact points, we first create a gripper mesh, transform it to the predicted pose $H$, and prune out any grasps that intersect with the object mesh. This is a relatively hard constraint as grasps that collide even slightly are pruned out. We then sample rays inwards from the gripper fingers and find the intersection with the ground truth mesh of the object, along with contact normals from the mesh faces. We then define the FC metric as the percentage of grasps that satisfy force closure out of the total number of predicted grasps (including the ones that collide).
@@
(A) Shape features without Convolutional Features: Convolutional plane features model local geometries efficiently by distilling features from existing pointcloud encoders like PointNet, as shown in Figure 2. Since we learn a joint representation for shapes and grasps (by converting them into points )[27], the features of a grasp pose $H$ are similar to the target region of the grasp. Thus we analyze how well the representation model’s the object geometry. To do this, we compare vanilla SE3Diff[28] with our model. Since both these models generate SDF as an auxiliary task, we analyze the role of convolutional features using Chamfer’s Distance (CD)  [^1] https://pdal.io/en/stable/apps/chamfer.html between pointclouds sampled from the predicted SDF and the ground truth pointcloud. As shown in Table II, our method significantly improves the reconstruction quality compared to SE3Diff, showcasing the effectiveness of Convolutional plane features.
(B) Part Conditioned Diffusion: It is possible to generate grasps on constraint regions with our method without using the part-guided strategy by employing a naive albeit sample-inefficient approach, i.e., by simply generating grasps on the target region pointcloud $P_{t}$ and thresholding grasps based on energy values (Equation 6). In this experiment, we try to see the effect our part-guided strategy has over this naive method in terms of sample efficiency. For this, we measure the percentage of grasps left after the thresholding step for grasps generated with and without the part-guided strategy (w/o PD). Since the bounds of energy are not fixed [29], we set $\delta$ such that the FC metric of the grasps from both the models is in a similar range as in Table I. As shown in Table II, the part-guided strategy increases the sample efficiency by almost 2.5 times.
@@ VI CONCLUSIONS
Future work could involve possible extensions to scenarios involving more than two arms, further improving the representation and formulating grasps with more variables, i.e. predicting the width, height, etc., potentially expanding CGDF’s application to a broader range of robotic manipulation tasks.

### v2

@@ IV CGDF: Constrained Grasp Diffusion Fields
Given a object point cloud $P\in\mathbb{R}^{N\times 3}$ along with a target region (i.e. constraint) $P_{t}\subseteq P$, our goal is to generate $M$ parallel-jaw grasp poses $H_{i}\in SE(3),i\in[0,M)$ on $P$, such that $H_{i}$ is located on $P_{t}$. Here $M$ is a number that can be user-defined. Additionally, we make no assumptions regarding the granularity of the target regions, allowing $P_{t}$ to encompass anything from a small, localized region to the entire point cloud. For an n-arm setting with large objects, $n$ target regions can be selected $P_{T}=\{P_{t}^{1},P_{t}^{2}\dots P_{t}^{n}\}$. Consequently, our method must learn a robust shape representation to generalize to arbitrary shapes and complex geometries, ensuring its applicability across various constraint regions. Further, the model should generate diverse grasps with full coverage over the target region. Diffusion models have shown excellent performance in generating diverse outputs while learning the data distribution effectively [28]. Thus we choose a diffusion-based architecture as our backbone.
@@ IV-A SE(3) Diffusion Model
$H_{k}=\text{Expmap}\left[\text{Logmap}(H)+\epsilon_{k}\right],\epsilon_{k}\sim\mathcal{N}(0,\sigma_{k}^{2}I)$ (3)
Loss Function: The diffusion model can now be trained to predict the noise given a perturbed sample by minimizing the L1 loss objective between the sampled perturbation $\epsilon_{k}$ and the predicted perturbation as follows:
@@
Inverse Diffusion Step: Following Se3diff [28], the inverse diffusion step is formulated as an adapted version of the Euclidean Langevin MCMC [32], as shown in Fig 2(c):
$H_{k-1}=\text{Expmap}\left(-\frac{\alpha_{k}^{2}}{2}s_{\theta}(H_{k},k,P)+\alpha_{k}\epsilon\right)H_{k}$ (5)
where $\epsilon\in\mathbb{R}^{6},\epsilon\sim\mathcal{N}(0,I)$ , $\alpha_{k}$ refers to a non-negative step-dependent coefficient.
@@ IV-B Constrained Grasp Diffusion
The above diffusion model trains on an unconstrained grasping objective and thus generates grasps covering the entire object pointcloud $P$. However, if the model is able to learn a rich enough shape representation to model local geometries, we can use the model trained on unconstrained grasping to generate constrained grasps without retraining, by using the energy values predicted by $E_{\theta}$ as explained below.
@@
Assume we have an object pointcloud $P$ and a target region pointcloud $P_{t}$. Since $E_{\theta}$ is an EBM, stable grasp poses have lower energies as compared to invalid or colliding grasps. Therefore, if we treat $P_{t}$ as a stand-alone pointcloud, then a grasp pose $H_{t}$ on it is stable if its energy at $k=0$ ($e^{\prime\prime}=E_{\theta}(H_{t},0,P_{t})$) is low (using Eqn 1). Now, we can also calculate the energy of the $H_{t}$ giving $P$ as the context instead of $P_{t}$, that is, $e^{\prime}=E_{\theta}(H_{t},0,P)$. Notice that $e^{\prime}$ is low only if $H_{t}$ is also a stable grasp on $P$. Thus, $H_{t}$ is a stable constrained grasp on $P$ if:
$e^{\prime}<\delta\And e^{\prime\prime}<\delta$
where $\delta$ is the user-defined energy threshold for valid grasps. This threshold depends on the training conditions of the model and can only be defined empirically by looking at the range of energy values returned by $E_{\theta}$.
@@
$s_{\theta}(H_{k},k,\{P,P_{t}\})=\frac{D\text{max}(e^{\prime}_{k},e^{\prime\prime}_{k})}{DH_{k}}$
@@
As shown in Equations 7 and 6, by taking the maximum value of the energies, we essentially guide the grasp from a random pose to a stable configuration near the target region. During the inverse diffusion step, if the grasp $H_{k}$ is near the constrained region but collides with the full object or is unstable, the energy $e^{\prime}_{k}$ is higher, which makes $H_{k}$ move to a more stable pose. Conversely, if the grasp is stable but is far from the constrained region, then $e^{\prime\prime}_{k}$ is higher, and the grasp moves closer to the constrained region. Eventually, the grasp moves to a pose where the energies for both $P$ and $P_{t}$ are low, i.e. the grasp is on the constrained region as well as stable.
@@ IV-C Model Architecture
[…]
@@
where $\oplus$ means concatenation. The grasp pose $H\in SE(3)$ is represented as $P_{H}=H\cdot P_{q}$, where $P_{q}\in\mathbb{R}^{N\times 3}$ is a fixed query pointcloud. The pose descriptors of $H$ can then be obtained as $f_{H}^{k}=\bigoplus_{i=0}^{N}[f_{x_{i}}^{k}],x_{i}\in P_{H}$. The decoder is an MLP that takes the grasp pose descriptors and outputs the energy of the grasp, $e_{k}=D_{\theta}(f_{H}^{k})$. The model trains on the main objective as defined in Equation 4, along with an auxiliary loss to predict the SDF as mentioned in [28].
@@ V Experiments
Datasets: We train and evaluate all models on grasps and objects from the DA2 Dataset  [6], which consists of dual-arm grasps sampled on arbitrary object meshes taken from  [34]. The meshes used in this datasets are taken from the ShapeNetSem dataset, which are scaled up relative to the grasp, making learning shape prior relatively challenging due to the large number of classes and fewer instances per class. The ACRONYM dataset  [15], also uses meshes from the same set but is catered towards smaller tabletop objects, and scales the large objects down to table-top/toy size, whereas DA2 keeps large objects like chairs, lamps, etc., in their original scale. Since the baseline models require a large-scale dataset consisting of object point clouds, along with successful grasps on randomly sampled targets, we augment the DA2222 dataset using the same procedure as described in  [11] Section V. We use the original DA2 dataset without augmentation to train our model. We further show qualitative results emphasizing the generalization capabilities of our model by combining our method with a language-based part segmentation model PartSLIP [5] on the PartNetE Dataset [35, 36]. As shown in Fig 1 and 4, we segment out the part using a text prompt (e.g.“handle of a microwave”). We then use our model with the part-guided diffusion strategy (trained on DA2) to generate grasps on the region.
@@
We evaluate the performance of these models using three metrics: Force Closure, Grasp Success Rate and Target Grasps. Force Closure serves as a stringent criterion, accounting for both stability and collisions between the object and the gripper in the generated dual-arm grasps. On the other hand, the Grasp Success Rate reflects a more practical assessment, simulating scenarios typical of physical interactions within an Isaac Gym environment [37]. Target Grasps measures the percentage of grasps that are generated near the target region.
@@
$GG^{\prime}\succcurlyeq\epsilon I_{6\times 6};\text{ and }\|Gc\|<\delta^{\prime}$ (8)
where $G$ is the grasp matrix,, $c$ is the axis of the friction cone, $S$ is the object surface and $x_{i}\in S$. Please refer to  [38] for more details. To get the contact points, we first create a gripper mesh, transform it to the predicted pose $H$, and prune out any grasps that intersect with the object mesh. This is a relatively hard constraint as grasps that collide even slightly are pruned out. We then sample rays inwards from the gripper fingers and find the intersection with the ground truth mesh of the object, along with contact normals from the mesh faces. We then define the FC metric as the percentage of grasps that satisfy force closure out of the total number of predicted grasps (including the ones that collide).
@@
(A) Convolutional Plane features: Convolutional plane features [12] represent local geometries efficiently by distilling features from existing pointcloud encoders like VN-PointNet, as shown in Figure 2. Since we learn a joint representation for shapes and grasps (by converting them into points )[27], the features of a grasp pose $H$ are similar to the target region of the grasp. Thus, we analyze how well this 3D representation encodes the object geometry. To do this, we compare vanilla SE3Diff[28] with our model. Since both these models generate SDF as an auxiliary task, we analyze the role of convolutional features using Chamfer’s Distance (CD)  [^1] https://pdal.io/en/stable/apps/chamfer.html between pointclouds sampled from the predicted SDF and the ground truth pointcloud. As shown in Table II,CGDF significantly improves the reconstruction quality compared to SE3Diff, showcasing the effectiveness of Convolutional plane features.
(B) Part Conditioned Diffusion: It is possible to generate grasps on constrained regions with CGDF without using the part-guided strategy, by employing a naive albeit sample-inefficient approach mentioned in Section IV-B, which is to simply generate grasps on the target region pointcloud $P_{t}$ and threshold grasps based on energy values. In this experiment, we try to study the effect of part-guided diffusion over this naive method in terms of sample efficiency. For this, we measure the percentage of grasps left after the thresholding step for grasps generated with and without part-guided diffusion (w/o PD). Since the bounds of energy are not fixed [29], we set the energy threshold $\delta$ such that the FC metric of the grasps from both the models is in a similar range as in Table I. As shown in Table II, part-guided diffusion generates 2.5 times times more valid grasps than the baseline (w/o PD).
@@ VI CONCLUSIONS
Future work could involve possible extensions to scenarios involving more than two arms, further improving the representation, potentially expanding CGDF’s application to a broader range of robotic manipulation tasks.